Please show me the work thank you
3) An object is thrown upward with a speed of 12 m/s on the surface of planet X where the acceleration due to gravity is 1.5 m/s2. How long does it take for the object to reach the maximum height? A) 8.0 s B) 11 s C) 14 s D) 16 s
Answer #1

ZaS v=0m/s s = (-112/2a)

acceleration is directed opposite to initial speed so we get negative in 'a'

so maximum height = 48 m

time taken = 12/1.5 = 8 sec

please rate it up thanks :)